OpenStudy (unofficialllyy):

Jose rented a truck for one day. There was a base fee of $14.99, and there was an additional charge of 96 cents for each mile driven. Jose had to pay $194.51 when he returned the truck. For how many miles did he drive the truck?

OpenStudy (unofficialllyy):

@jim_thompson5910

OpenStudy (unofficialllyy):

@jdoe0001

jimthompson5910 (jim_thompson5910):

were you able to figure out the total cost equation?

OpenStudy (unofficialllyy):

No i'm really bad at word problems... I get confused on what goes where.

jimthompson5910 (jim_thompson5910):

Start with `There was a base fee of $14.99`

jimthompson5910 (jim_thompson5910):

then add on `additional charge of 96 cents for each mile driven` so if x = number of miles driven, then it costs an additional 0.96x dollars to drive x miles the total cost equation would be y = 14.99 + 0.96x agreed?

OpenStudy (unofficialllyy):

Yes

jimthompson5910 (jim_thompson5910):

`Jose had to pay $194.51 when he returned the truck` so his total cost is y = 194.51

jimthompson5910 (jim_thompson5910):

y = 14.99 + 0.96x 194.51 = 14.99 + 0.96x ... replace y with 194.51 solve for x

OpenStudy (unofficialllyy):

187 miles

jimthompson5910 (jim_thompson5910):

correct